Circular No. 7412

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

COMET C/2000 H2 (SOHO) D. A. Biesecker, SM&A Corporation and Goddard Space Flight Cen-

ter, reports observations of another Kreutz sungrazing comet found by T. Lovejoy, M. Boschat, M. Oates, and R. Gorelli via the SOHO website. The comet was observed in both the C2 and C3 SOHO telescopes, and a tail at least 0^{o}.5 long was visible. The comet evidently became brighter than mag 1, but at peak brightness, the coma saturated the CCD detector. Astrometry (measured by D. Hammer and Biesecker, reduced by B. G. Marsden) and orbital elements appear on MPEC 2000-J05.

2000 UT  α_{2000}  δ_{2000}
Apr. 28.513  2^{h}45.^{m}2  +11^{o}21′

XTE J1118+480 J. Patterson, Columbia University, reports on behalf of the Center for

Backyard Astrophysics (CBA) project that 27 nights of photometric observation of the optical counterpart of the x-ray transient XTE J1118+480 by CBA observers (including R. Fried, D. Harvey, L. Cook, G. Masi, J. Kemp, J. Hannon, D. Buczynski, T. Vanmunster, L. Pidgley, P. Maxted, and S. Araujo-Betancor) yield a precise period of 0.17082(11) day and full amplitude 0.052 mag. Patterson adds: “Variability on other (shorter and longer) timescales is largely absent, except at very short periods. Largeamplitude (> 0.3 mag) flickering is present on timescales of < 10 s, in all passbands (U through I). Observations with very high time resolution are needed to define this remarkable phenomenon. More details are available

at http://www.astro.bio2.edu/cba.”

SUPERNOVA 2000bs IN UGC 10710 D. Balam, University of Victoria, reports the following Johnson B and

V and Cousins R and I photometry obtained with the 1.82-m Plaskett telescope of the National Research Council of Canada on Apr. 24.48 UT: V = 17.37 ± 0.02, B − V = +0.63 ± 0.04, V − R = +0.31 ± 0.03, R − I = +0.27 ± 0.04.

Additional R CCD magnitude estimates (uncertainty ± 0.2 mag) by K. Hornoch, Lelekovice, Czech Republic (0.35-m reflector): Apr. 23.960 UT, 17.2; 29.009, 17.3; 29.889, 17.2; May 1.000, 17.3.

2000 May 1  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2000 CBAT  Daniel W. E. Green